The length of a rectangle is 5 in longer than its width.
If the perimeter of the rectangle is 42 in, find its area.

Answer by TimothyLamb(4379)

---
x = y + 5
2x + 2y = 42
---
put the system of linear equations into standard form:
---
x - y = 5
2x + 2y = 42
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 13 in
y = 8 in
area = 8*13 = 104 sq.in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php